OpenStudy (anonymous):

The integral of cos(u^2)du from 5x to cos(x). Not looking for an answer, looking for steps.

OpenStudy (amistre64):

can we sub z^(1/2) for x ?

OpenStudy (anonymous):

Dunno.

OpenStudy (amistre64):

err.... i meant sub it for u lol

OpenStudy (anonymous):

Then what do we get cos(x^4)?

OpenStudy (amistre64):

if u = z^(1/2) then du = \(\frac{1}{2z^{1/2}}\) dz \[\int\frac{cos(z)}{2z^{1/2}}dz\]

OpenStudy (amistre64):

why does the latex z look like an x ?

OpenStudy (anonymous):

That looks like it just made the problem worse.

OpenStudy (anonymous):

Nevermind, Z for X

OpenStudy (anonymous):

Do I have to put the limits in terms of x even though they have x's in them?

OpenStudy (amistre64):

lol .... sometimes it gets better sometimes not

OpenStudy (amistre64):

the limits arent the issue are they? if we can integrate the cos(u^2) the limits are a sinch

OpenStudy (anonymous):

Well when you use U sub with limits you usually change them right?

OpenStudy (amistre64):

table of ints says: cos(ax^2) goes to: sqrt(pi/2a) C(sqrt{a}); but i got no idea what the C means

OpenStudy (amistre64):

i always mess it up when I change limits, so I just convert back to the original in the end

OpenStudy (anonymous):

So I don't need to change the limits, I only needed to put the integral itself into terms of x?

OpenStudy (amistre64):

the variable of the integral just tells you what to do to the limits after integrating; spose it ints up to sin(u); then whenever you see a u, replace it with the limit to get: sin(cos(x)) - sin(5x)

OpenStudy (amistre64):

wolfram gives me the same fresnel integral

OpenStudy (anonymous):

Answer is Sin(x)*cos(cos^2x)+5(cos(25x^2)) @_@

OpenStudy (amistre64):

is that from an answer sheet?

OpenStudy (anonymous):

Yea

OpenStudy (amistre64):

i cant see how its done ... :)